OpenStudy (czesc):

Cartesian equation for the polar equation r sin θ = 6?

jimthompson5910 (jim_thompson5910):

Make sure these two equations are somewhere in your notes or on flashcards \[\Large x = r*\cos(\theta)\] \[\Large y = r*\sin(\theta)\] they are very useful for converting from polar to cartesian form

OpenStudy (czesc):

so y=6? @jim_thompson5910

OpenStudy (usukidoll):

I also got y =6 when I converted as well, but...it's been a while since I've done this

jimthompson5910 (jim_thompson5910):

yep, this polar equation is a horizontal line through y = 6

OpenStudy (usukidoll):

ah yea...